OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4

OpenStudy (anonymous):

Typo!! y less than or equal to x - 5 y greater than or equal to -x - 4

OpenStudy (inkyvoyd):

We can't actually do anything without the graph.

OpenStudy (apoorvk):

So the two equations are: y <= x-5 and y>=-x-4 Now since you haven't posted which "following points". Just substitute every point in both equation and verify if it satisfies conditions imposed by BOTH the equations. The one doing so would be the answer. Or, the graphing method as inky says. Which is better.

OpenStudy (anonymous):

Draw a graph for both lines y=x-5 and y=-x-4 then identify the region where y<=x-5 and y>=-x-4. the intersection of both regions is the answer.